OpenStudy (anonymous):

Ok so how would I work this out? Please do not give me the answer just explain to me how I can solve it thnxs Write each answer as a power of 2. Show your work and explain your steps. a. Computer capacity is often measured in bits and bytes. A bit is the smallest unit, which is a 1 or 0, in the computer's memory. A byte is 2^3 bits. A megabyte (MB) is 2^20 bytes. How many bits are in a megabyte?

OpenStudy (anonymous):

well if i said a bit is 8 and a megabit is 24 how many bits to a megabyte ? 23/8 therefore 3 bits. Try the same logic here

OpenStudy (anonymous):

24/8 my bad

OpenStudy (anonymous):

8 of what and 24 of what?

OpenStudy (anonymous):

just anything really, i chose arbitrary numbers to show a concept. But i did explain it badly Basically you have to look at the question as asking. How many Bits go into A byte which you are told is 2^3 Then you know there are 2^20 bytes to a megabyte so to find the amount of bits, convert the 2^20 bytes to (this is what your looking for) amount of bits

OpenStudy (akashdeepdeb):

1byte = 2^3 bits 1 megabyte = 2^20 bytes! So 1 megabyte = 2^20 bytes OR 2^20 * 2^3 BITS = 2^23 bits Got it? :)

OpenStudy (akashdeepdeb):

@Taplin44 : Has done it in a better way though.

OpenStudy (anonymous):

No :'(

OpenStudy (akashdeepdeb):

1 kilometer is 1000 metres! 1 metre is 100centimeters How many centimeters is 1 kilometers?

OpenStudy (anonymous):

One thing I don't under stand is it says o write it in a power of 2 and also it say a bit is 0 or 1

OpenStudy (akashdeepdeb):

This is a bit = 0 or 1 This maybe a byte: 01000101

OpenStudy (anonymous):

This is confusing me...

OpenStudy (anonymous):

10,000 centimeters

OpenStudy (akashdeepdeb):

1 miunte is 60 seconds 1 second is 100 miliseconds How many miliseconds is 1 minute!

OpenStudy (anonymous):

dont get bogged on the complications of the question, just identify what it is asking. When it says as a power of two it means do the calculation AkashdeepDeb showed 1 megabyte = 2^20 bytes OR (2^20) * (2^3) BITS =2^23 (see how it is 2^something) what a byte actually is is meaningless here

OpenStudy (akashdeepdeb):

It actually is 100,000 centimeters 100 * 1000 Getting it?

OpenStudy (anonymous):

Ok but how do I do the problem I still don't under stand

OpenStudy (akashdeepdeb):

1 miunte is 60 seconds 1 second is 100 miliseconds How many miliseconds is 1 minute!

OpenStudy (anonymous):

6000?

OpenStudy (akashdeepdeb):

Excellent!! Now that IS what the question asks you! How many bits are in a megabyte? 1 byte = 2^3 bits 1 megabyte = 2^20 bytes 1megabyte = ? bits = 2^20 * 2^3

OpenStudy (akashdeepdeb):

That IS how you solved the other questions right? Solve this in the same way!! :D

OpenStudy (anonymous):

Ok so 1 megabyte = 2^23 Bits?

OpenStudy (akashdeepdeb):

YES!! :D

OpenStudy (akashdeepdeb):

Understood? :)

OpenStudy (anonymous):

Yes but do I need to work out the exponent for this question

OpenStudy (akashdeepdeb):

Exponent = 23! How? Like this.... 2^20 * 2^3 = 2^(20+3) = 2^23 bits Got it? :)

OpenStudy (anonymous):

Yes I got that but I mean like do I need to work it out like do 2*2*2*2*2 & so on for this?

OpenStudy (akashdeepdeb):

No No Noo!!! No need!! :D Hahah.

OpenStudy (anonymous):

Ok good lol thanks

OpenStudy (akashdeepdeb):

:)

OpenStudy (anonymous):

Here is the other part to the question I need some help on it to if you could please help A gigabyte (GB) is 2^30 bytes. How many Megabytes are in a gigabyte? How many bits are in a gigabyte? Ok so would it be 1 Mega bytes with 2^7 left over?

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (akashdeepdeb):

1 gb = 2^30 bytes 1 mb = 2^20 bytes 1gb = 2^30/2^20 = 2^10 mb And for 1 gb to bits 1 gb = 2^30 * 2^3 bits [Because 1 bytes is 2^3 bits] = 2^33 bits.

OpenStudy (akashdeepdeb):

I am a little bit busy now. Whenever I am jobless I'll show you a cool method called the unit factor method. THat'll help you a lot in conversions.

OpenStudy (anonymous):

Ok thanks

OpenStudy (akashdeepdeb):

:)